OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 64 cm3? = ? cm

OpenStudy (anonymous):

would i do 64 divided by 6 cause a cube has 6 sides?

OpenStudy (arbershabani97):

x^3=64, just see another example so you will know how to do it ( i will choose 125) x^3=125 x^3=5^3 x=5

OpenStudy (isaiah.feynman):



OpenStudy (anonymous):

and i think the answer is 32....cause it says 64 cm3?

OpenStudy (anonymous):

64x3=192 and 192 divided by 6 is 32

OpenStudy (arbershabani97):

64=x^3 4^3=x^3 (x*x*x=64)

OpenStudy (isaiah.feynman):

It can't be 32.